OpenStudy (anonymous):

Two angles of a triangle have the same measure and the third one is 33 degrees greater than the measure of each of the other two. Find the measure of the LARGEST angle in the triangle. Answer: The LARGEST angle has a measure of

OpenStudy (anonymous):

 Total angle in a triangle=180 (33+x)+x+x=180 Solve for x Then plug it in here: 33+x

OpenStudy (dn_20):

We have: x +x + x+33 = 180 3x +33 = 180 3x = 147 x = 49 plug 49 for x +33 we have: x +33 = 49 + 33 = 82

OpenStudy (dn_20):

So the largest angle measure are 82